### v1

# Path Planning in Complex Environments with Superquadrics and Voronoi-Based Orientation

Abstract
Path planning in narrow passages is a challenging problem in various applications. Traditional planning algorithms often face challenges in complex environments like mazes and traps, where narrow entrances require special orientation control for successful navigation. In this work, we present a novel approach that combines superquadrics (SQ) representation and Voronoi diagrams to solve the narrow passage problem in both 2D and 3D environment. Our method utilizes the SQ formulation to expand obstacles, eliminating impassable passages, while Voronoi hyperplane ensures maximum clearance path. Additionally, the hyperplane provides a natural reference for robot orientation, aligning its long axis with the passage direction. We validate our framework through a 2D object retrieval task and 3D drone simulation, demonstrating that our approach outperforms classical planners and a cutting-edge drone planner by ensuring passable trajectories with maximum clearance.
Index Terms:
Voronoi diagrams, Superquadrics, Motion planning, Narrow passage.
I Introduction
Path planning in narrow passages poses challenges across various applications, including robotic manipulation [1, 2], mobile robotics [3], and drone navigation [4]. The task involves guiding a robot to navigate through a narrow passage without detouring or colliding with obstacles. The challenge arises from the fact that not all robots or obstacles can be approximated as simple geometric shapes, such as spheres or grid, though this is a common simplification in mobile robots. This approximation allows for easier handling of collision detection and path planning algorithms, as it reduces the complexity of the vehicle’s geometry. However, in many cases, robots must find a path through the passage while maintaining a correct orientation. For example, as depicted in Fig. 1(b), a manipulator attempts to retrieve an object from a crowded shelf. Due to height constraints, the robot is restricted to navigating in a 2D plane, as shown in Fig. 1(a). Some passages may be too narrow for the robot to traverse, while others are wide enough to accommodate the object’s shorter axis length, providing several potential trajectories, shown in green. However, due to perception uncertainties or control inaccuracies, some of these trajectories may result in collisions with obstacles if they are too close to them (Fig.1(c)). Similarly, in drone navigation, certain passages may require the drone to tilt at specific angles to pass through, while in other cases, the robot must recognize that certain passages are impassable, as illustrated in Fig.1(d).
(a) Planning in deterministic case: identifying passable and impassable routes
(b) Object retrieval in cluttered spaces: cross-sectional view
(c) Uncertainty in real-world: potential collisions due to perception errors
(d) 3D Navigation through narrow passages: a drone pass through narrow pillars
Figure 1: Navigating narrow passages in 2D and 3D: path feasibility, orientation decision, and safety problem due to uncertainty
Many traditional path planning algorithms are popular and widely applied due to their favorable performance. However, some fall short when dealing with complex environments like mazes, traps, and narrow passages. For example, while sampling-based planning methods perform well in high-dimensional tasks [5], they struggle with efficiently sampling and growing a path through narrow passages. Similarly, artificial potential field methods are widely used for obstacle avoidance [6], but they are prone to getting stuck in local optima. One effective approach for navigating narrow passages and mazes is the Voronoi diagram [7, 8, 9]. This method provides the safest solution for the planner by ensuring maximum clearance [10], which means the robot remains as far as possible from all obstacles. Additionally, some algorithms have already achieved near real-time efficiency in mobile robot path planning [11, 12] with the help of Voronoi diagram. However, these approaches lack pose control in narrow passage scenarios because circular mobile robots typically do not need orientation control in such environments.
The choice to approximate objects is crucial. A model that is too simplified (e.g., a sphere) may prevent the robot from passing through certain narrow passages. For instance, if the green ellipse in Fig. 1(a) is approximated as a sphere, the robot would lose the opportunity to retrieve the object. On the other hand, using an overly accurate model can result in computational overhead during collision detection. Orthey et al. [13] successfully tackled narrow passage problems by employing an efficient yet complex approach. Their method, while powerful, requires numerous patterns and a detailed CAD model of the object for collision detection.
In contrast, we prioritize a balance between simplification and efficiency by representing objects using superquadrics (SQ) [14], which provide both differentiability and a convenient parametrization. SQ can describe a wide range of geometric shapes. For example, Ruan et al. [15] address narrow passage problems using SQ to represent both robots and obstacles, employing Minkowski sums to avoid collisions across various robot orientations. Their approach achieved effective planning for self-collision avoidance. Zhao et al. [16] simplify the Minkowski sum via utilizing the formulation of SQ. For now, our task is simpler than Ruan et al. [15] and thus, we do not require extensive computation with Minkowski sums. Additionally, we avoid over-simplification, as many researchers use Minkowski sums to grow obstacle sizes by sweeping the maximum size of the robot, ensuring the remaining map is navigable with any robot pose.
In this work, we propose a simpler and faster approach for addressing the narrow passage problem in both 2D and 3D scenarios, combining SQ and Voronoi diagrams. Our method leverages two key advantages of SQs: first, we expand obstacles using the SQ size parameter by adding the robot’s minimum axis length, effectively eliminating impassable passages. Second, by parameterizing SQs in polar coordinates, we can easily compute hyperplanes between them and generate a Voronoi diagram. This ensures a maximum clearance path, and the hyperplanes naturally guide the robot’s orientation, aligning its long axis with the passage. Our framework is applicable to both 2D and 3D scenarios. We validate it through a 2D object retrieval task, successfully identifying the correct passage and orientation, and we compare it with classical planners. Additionally, simulations in 3D show that our method outperforms a cutting edge drone planner by finding passable paths while maintaining maximum clearance.
II Preliminaries: Superquadrics and Voronoi diagram
(a) Original obstacles.
(b) Grown obstacles.
(c) Hyperplane.
(d) Create Voronoi diagram.
(e) planning via DMP.
Figure 2: An entire process of our algorithm: a): approximate the environment by SQs. b): Grown obstacles with shortest axis length of robot. c): Hyperplane between two SQs and closest point. d): Create Voronoi diagram with hyperplane. e): Utilize graph planning to find edges, then parameterized by DMP.
In this section, we present the mathematical definitions of Superquadrics (SQ) (known as Superellipses in 2D) and the Voronoi diagram. Superquadrics form a family of three-dimensional shapes that generalize basic quadric surfaces, such as ellipsoids, elliptic paraboloids, and hyperbolic paraboloids. These surfaces are generated by the spherical product of two parametric two-dimensional curves [14].
II-A Superellipses and superquadrics
Superquadrics are widely utilized to represent geometry of objects in real world [17]. The 2D analogue of superquadrics are superellipses,
II-A1 Superellipses
Superellipses is implicitly defined via equation:
$\left(\frac{x}{a_{1}}\right)^{\frac{2}{\epsilon}}+\left(\frac{y}{a_{2}}\right)^{\frac{2}{\epsilon}}=1$ (1)
Here, $\epsilon$ represents the exponent of the superellipse, which controls the shape of the superellipse. $a_{1},a_{2}$ control the size. Meanwhile, define a inside-outside function $F(x,y)$ as
$F(x,y)$ $=\left(\frac{x}{a_{1}}\right)^{\frac{2}{\epsilon}}+\left(\frac{y}{a_{2}}\right)^{\frac{2}{\epsilon}}-1$ (2)
The function $F(x,y)$ possesses a very useful property. For a point $(x_{0},y_{0})$, if $F(x_{0},y_{0})>0$, then it is outside the SQ, if $F(x_{0},y_{0})=0$ then it is on the surface, and if $F(x_{0},y_{0})<0$, then it is inside. For this reason, the function F is also referred to as the inside-outside function.
Meanwhile, analogous to a circle, superellipse can be parameterized as [14],
$\mathbf{r}(\omega)$ $=\begin{bmatrix}a_{1}\cos^{\epsilon}\omega\\ a_{2}\sin^{\epsilon}\omega\end{bmatrix},$ (3a)
$\omega\in[-\pi,\pi]$ (3b)
The exponentiation with power $\epsilon$ here is the signed power function, i.e
$\cos^{\epsilon}\omega=\operatorname{sign}(\cos\omega)\cdot\left|\cos\omega\right|^{\epsilon}$ (4)
II-A2 Superquadrics
In 3D case, the implicit function of superquadrics is defined as,
$\left(\left(\frac{x}{a_{1}}\right)^{2/\epsilon_{2}}+\left(\frac{y}{a_{2}}\right)^{2/\epsilon_{2}}\right)^{\epsilon_{2}/\epsilon_{1}}+\left(\frac{z}{a_{3}}\right)^{2/\epsilon_{1}}$ $=1$ (5)
So we have a similar inside-outside function $F(x,y,z)$ (analgoy to Eq. 2 for 3D case)
$F(x,y,z)$ $=\left(\left(\frac{x}{a_{1}}\right)^{2/\epsilon_{2}}+\left(\frac{y}{a_{2}}\right)^{2/\epsilon_{2}}\right)^{\epsilon_{2}/\epsilon_{1}}+\left(\frac{z}{a_{3}}\right)^{2/\epsilon_{1}}-1$ (6)
$F(x,y,z)$ is still inside-outside function in 3D case. Meanwhile, Eq. 5 can be parameterized by
$\mathbf{r}(\eta,\omega)$ $=\begin{bmatrix}a_{1}cos^{\epsilon_{1}}\eta~{}cos^{\epsilon_{2}}\omega\\ a_{2}cos^{\epsilon_{1}}\eta~{}sin^{\epsilon_{2}}\omega\\ a_{3}sin^{\epsilon_{1}}\eta\end{bmatrix},$ (7a)
$\eta\in[-\pi/2,\pi/2],~{}\omega\in[-\pi,\pi]$ (7b)
Here, $(a_{1},a_{2},a_{3})$ are the size parameters, while $(\epsilon_{1},\epsilon_{2})$ are the shape parameters. Thus, in local coordinates of the superquadric, 5 parameters are needed to define it completely.
II-A3 Superellipses and Superquadrics in general position
To accommodate any orientation and position, translation and rotation operations must be applied to the superellipses and superquadrics. Transformation matrices are required to represent superellipses and superquadrics in the world frame. Therefore, to fully describe a SQ, the parameter set is defined as $SQ:=[{\boldsymbol{\mathbf{\epsilon}}},{\boldsymbol{\mathbf{a}}},{\boldsymbol{\mathbf{p}}},{\boldsymbol{\mathbf{\theta}}}]^{T}$, where ${\boldsymbol{\mathbf{T}}}({\boldsymbol{\mathbf{p}}},{\boldsymbol{\mathbf{\theta}}})=[{\boldsymbol{\mathbf{R}}}({\boldsymbol{\mathbf{\theta}}})\in SO(3),{\boldsymbol{\mathbf{p}}}\in\mathbb{R}^{3}]\in SE(3)$ is the transformation matrix to describe the pose of SQ. For any point ${}^{\mathcal{S}}{\boldsymbol{\mathbf{c}}}$ in SQ’s local coordinates $\mathcal{S}$, its corresponding position in the world coordinate system is given by:
${\boldsymbol{\mathbf{c}}}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}{\boldsymbol{\mathbf{T}}}_{\mathcal{S}}{{}^{\mathcal{S}}\tilde{{\boldsymbol{\mathbf{c}}}}}$ (8)
where
${}^{\mathcal{S}}\tilde{{\boldsymbol{\mathbf{c}}}}=\begin{bmatrix}{}^{\mathcal{S}}x_{c}\\ {}^{\mathcal{S}}y_{c}\\ {}^{\mathcal{S}}z_{c}\\ 1\end{bmatrix}$ (9)
II-B Voronoi Diagrams
Fundamental data structures play a pivotal role in computational geometry, enabling efficient algorithms for solving various geometric problems. One such essential structure is the Voronoi diagram, which partitions the Euclidean space $W\in\mathbb{R}^{n}$ into $m$ half spaces $\nu=\{V_{1},V_{2},\ldots,V_{m}\}$, often referred to as cells or regions,
Each cell is defined such that every point within it is closest to the corresponding disjoint convex obstacles $\mathcal{O}=\{O_{1},O_{2},\ldots,O_{m}\}$ (i.e., $d(O_{i},O_{j})>0\ \forall i\neq j$), based on maximum margin separating hyperplane, to be
$V_{i}$ :=$:=\left\{q\in W\ \middle|\ \|q-p_{i}^{*}\|\leq\|q-p_{j}^{*}\|\right\},$
$(p_{i}^{*},p_{j}^{*})$ $=\underset{p_{i}\in\mathcal{O}_{i},p_{j}\in\mathcal{O}_{j}}{\arg\min}d(p_{i},p_{j})\quad\forall j\neq i$ (10)
This can also be proved by separating hyperplane theorem [18] For any two non-intersecting convex sets $A,B\in\mathbb{R}^{n}$ (i.e., $A\cap B=\emptyset$), there exists $a\in\mathbb{R}^{n}$ and $b\in\mathbb{R}$ such that $a^{T}x\geq b$ for all $x\in A$ and $a^{T}x\leq b$ for all $x\in B$. This partitioning ensures that each region captures the space surrounding a specific object i.e., $O_{i}\subset V_{i}$, thereby facilitating spatial analysis and problem-solving within computational geometry.
Standard Voronoi diagram has point obstacles, say $O_{i}=\{p_{i}\}$ for some $p_{i}\in\mathbb{R}^{n}$, the generalized Voronoi diagram of $W$ in Eq.10 reduces to the standard Voronoi diagram of $W$, generated by points $\{p_{1},\ldots,p_{m}\}$, i.e., $V_{i}=\left\{q\in W\ \middle|\ \|q-p_{i}\|\leq\|q-p_{j}\|\ \forall j\neq i\right\}$ [19].
III robot navigation in generalized world
Many objects in the real world can be approximated, or decomposed and then approximated, by one or several SQs [17]. Unlike many approaches that assume all obstacles are disjoint and that the distance between obstacles is greater than the robot’s maximum diameter, our SQ-based representation allows for a more generalized and realistic world.
III-A Eliminate impossible passage by expanding
In robot planning, mobile robots [12] and drones [20] are commonly approximated by a sphere, after which collision checks or Minkowski sums are applied to all obstacles.
However, this method restricts the ability of non-spherical robots, such as a truck or a peg-like object, from passing through narrow passages. Therefore, in our approach, the first step is to approximate the robot, object, or drone using a Superquadric, represented as $SQ_{r}=[{\boldsymbol{\mathbf{\epsilon}}}_{r},{\boldsymbol{\mathbf{a}}}_{r},{\boldsymbol{\mathbf{p}}}_{r},{\boldsymbol{\mathbf{\theta}}}_{r}]^{T}$, while the other obstacles are represented as $SQ_{o}=[{\boldsymbol{\mathbf{\epsilon}}}_{o},{\boldsymbol{\mathbf{a}}}_{o},{\boldsymbol{\mathbf{p}}}_{o},{\boldsymbol{\mathbf{\theta}}}_{o}]^{T}$. For example, approximate some obstacles like Fig. 2(a).
We arrange the components of $a_{i}$ in ascending order, denoted as ${\boldsymbol{\mathbf{a}}}=[a_{1},a_{2},a_{3}],a_{i}<a_{i+1}$ so $a_{r,1}$ is the shortest axis of $SQ_{r}$. We expand this shortest axis to evaluate whether the robot can navigate through a narrow passage.
While the original concept of Voronoi-based planning is effective for point robots, it cannot be directly applied to robots with extended bodies. The Minkowski sum, a fundamental operation in computational geometry and robotics, is commonly used to handle this situation. Given two sets $A$ and $B$ in a vector space, the Minkowski sum $A+B$ is defined as the set of all points that can be expressed as the sum of a point from $A$ and a point from $B$: $A+B=\{a+b\ |\ a\in A,b\in B\}$. Instead of applying the Minkowski sum directly, we leverage the properties of Superquadrics to achieve similar results by expanding ${\boldsymbol{\mathbf{a}}}_{o}$ to achieve similar performance of Minkowski sum. This expansion yields an adjusted obstacle, represented $SQ^{M}_{o}$, where the size parameters are modified accordingly to reflect the expanded configuration.
${\boldsymbol{\mathbf{a}}}^{M}$ $={\boldsymbol{\mathbf{a}}}+a_{r,1}{\boldsymbol{\mathbf{I}}}$ (11)
where ${\boldsymbol{\mathbf{I}}}$ denotes the identity matrix. Now the robot is reduced to a point, as shown in Fig. 2(b), on which we can apply standard voronoi diagram.
III-B Generate Voronoi diagram
One advantage of using SQ is that they can be parameterized by angular coordinates, making them continuous functions. This property allows us to efficiently compute the minimum distance between SQ pairs using classical optimization algorithms to identify the closest points between each SQ pair. As illustrated in Fig. 2(c), two obstacles are represented by two SQs. The red circles indicate the points $(p_{i}^{*},p_{j}^{*})$, where the distance $d(p_{i},p_{j})$ is minimized, and the green plane represents the hyperplane that separates the two regions.
Meanwhile, we evaluate whether $F_{i}(p_{j}^{*})\leq 0$. If this condition holds, the two SQs overlap, and we assign them to the same cluster $C_{i}$. Otherwise, we separate two SQ as two clusters $C_{i}$ and $C_{j}$. The logic is implemented in Alg. 1.
Algorithm 1 Generate cluster
1: the obstacles after mink sum ${\boldsymbol{\mathbf{S}}}Q^{M}_{o}$, the world space $W$.
2: the obstacles after mink sum ${\boldsymbol{\mathbf{S}}}Q^{M}_{o}$, the world space $W$. the obstacles after mink sum ${\boldsymbol{\mathbf{S}}}Q^{M}_{o}$, the world space $W$. procedure Create Cluster(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$) the obstacles after mink sum ${\boldsymbol{\mathbf{S}}}Q^{M}_{o}$, the world space $W$. the obstacles after mink sum ${\boldsymbol{\mathbf{S}}}Q^{M}_{o}$, the world space $W$. procedure Create Cluster(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$)
3:      the obstacles after mink sum ${\boldsymbol{\mathbf{S}}}Q^{M}_{o}$, the world space $W$. the obstacles after mink sum ${\boldsymbol{\mathbf{S}}}Q^{M}_{o}$, the world space $W$. procedure Create Cluster(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$) Cluster $C\leftarrow\emptyset$, cluster index $id=1$. Cluster $C\leftarrow\emptyset$, cluster index $id=1$.
4:      Cluster $C\leftarrow\emptyset$, cluster index $id=1$. Cluster $C\leftarrow\emptyset$, cluster index $id=1$. Cluster $C\leftarrow\emptyset$, cluster index $id=1$. for $i=1,\dots,N$ do Cluster $C\leftarrow\emptyset$, cluster index $id=1$. Cluster $C\leftarrow\emptyset$, cluster index $id=1$. for $i=1,\dots,N$ do
5:          Cluster $C\leftarrow\emptyset$, cluster index $id=1$. Cluster $C\leftarrow\emptyset$, cluster index $id=1$. for $i=1,\dots,N$ do Cluster $C\leftarrow\emptyset$, cluster index $id=1$. Cluster $C\leftarrow\emptyset$, cluster index $id=1$. for $i=1,\dots,N$ do Cluster $C\leftarrow\emptyset$, cluster index $id=1$. Cluster $C\leftarrow\emptyset$, cluster index $id=1$. for $i=1,\dots,N$ do if $SQ^{M}_{o,i}\notin C$ then Cluster $C\leftarrow\emptyset$, cluster index $id=1$. Cluster $C\leftarrow\emptyset$, cluster index $id=1$. for $i=1,\dots,N$ do Cluster $C\leftarrow\emptyset$, cluster index $id=1$. Cluster $C\leftarrow\emptyset$, cluster index $id=1$. for $i=1,\dots,N$ do if $SQ^{M}_{o,i}\notin C$ then
6:               Cluster $C\leftarrow\emptyset$, cluster index $id=1$. Cluster $C\leftarrow\emptyset$, cluster index $id=1$. for $i=1,\dots,N$ do Cluster $C\leftarrow\emptyset$, cluster index $id=1$. Cluster $C\leftarrow\emptyset$, cluster index $id=1$. for $i=1,\dots,N$ do if $SQ^{M}_{o,i}\notin C$ then $id\leftarrow id+1$ $id\leftarrow id+1$
7:               $id\leftarrow id+1$ $C_{id}\leftarrow\emptyset$ $C_{id}\leftarrow\emptyset$ $C_{id}\leftarrow\emptyset$
8:          $C_{id}\leftarrow\emptyset$ $C_{id}\leftarrow\emptyset$ $C_{id}\leftarrow\emptyset$ for $j=i,\dots,N$ do $C_{id}\leftarrow\emptyset$ $C_{id}\leftarrow\emptyset$ for $j=i,\dots,N$ do
9:               $C_{id}\leftarrow\emptyset$ $C_{id}\leftarrow\emptyset$ for $j=i,\dots,N$ do $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{p_{i}\in SQ^{M}_{o,i},p_{,}j\in SQ^{M}_{o,j}}{\arg\min}d(p_{i},p_{j})$, $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{p_{i}\in SQ^{M}_{o,i},p_{,}j\in SQ^{M}_{o,j}}{\arg\min}d(p_{i},p_{j})$,
10:               $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{p_{i}\in SQ^{M}_{o,i},p_{,}j\in SQ^{M}_{o,j}}{\arg\min}d(p_{i},p_{j})$, $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{p_{i}\in SQ^{M}_{o,i},p_{,}j\in SQ^{M}_{o,j}}{\arg\min}d(p_{i},p_{j})$, $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{p_{i}\in SQ^{M}_{o,i},p_{,}j\in SQ^{M}_{o,j}}{\arg\min}d(p_{i},p_{j})$, if $F_{i}(p_{j}^{*})\leq 0$ then $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{p_{i}\in SQ^{M}_{o,i},p_{,}j\in SQ^{M}_{o,j}}{\arg\min}d(p_{i},p_{j})$, $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{p_{i}\in SQ^{M}_{o,i},p_{,}j\in SQ^{M}_{o,j}}{\arg\min}d(p_{i},p_{j})$, if $F_{i}(p_{j}^{*})\leq 0$ then
11:                   $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{p_{i}\in SQ^{M}_{o,i},p_{,}j\in SQ^{M}_{o,j}}{\arg\min}d(p_{i},p_{j})$, $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{p_{i}\in SQ^{M}_{o,i},p_{,}j\in SQ^{M}_{o,j}}{\arg\min}d(p_{i},p_{j})$, if $F_{i}(p_{j}^{*})\leq 0$ then $C_{i}\leftarrow C_{i}\cup SQ^{M}_{o,i}$ $C_{i}\leftarrow C_{i}\cup SQ^{M}_{o,i}$ $C_{i}\leftarrow C_{i}\cup SQ^{M}_{o,i}$
12:          $C_{i}\leftarrow C_{i}\cup SQ^{M}_{o,i}$ $C\leftarrow C\cup C_{i}$ $C\leftarrow C\cup C_{i}$ $C\leftarrow C\cup C_{i}$
13:      $C\leftarrow C\cup C_{i}$ return $C$ return $C$ return $C$
After separating the clusters based on the overlapping of SQs, we proceed to create the diagram. Similarly, we loop through all the clusters. For any the hyperplane $HP_{i,j}$ between SQ $i$ and SQ $j$ can be expressed at the center of the $\frac{p^{*}_{i}+p^{*}_{j}}{2}$. Meanwhile, define the normalize function $\hat{n}(x)$ as
$\hat{n}({\boldsymbol{\mathbf{x}}})=\frac{{\boldsymbol{\mathbf{x}}}}{\norm{{\boldsymbol{\mathbf{x}}}}}$ (12)
Thus, the normal vector of $HP_{i,j}$ equals,
${\boldsymbol{\mathbf{n}}}_{i,j}=\hat{n}({\boldsymbol{\mathbf{p}}}^{*}_{i}-{\boldsymbol{\mathbf{p}}}^{*}_{j})$ (13)
Following the definition of hyperplane, we get
$HP_{i,j}=\{x|{\boldsymbol{\mathbf{n}}}_{i,j}^{T}x\leq{\boldsymbol{\mathbf{n}}}_{i,j}^{T}(\frac{{\boldsymbol{\mathbf{p}}}^{*}_{i}+{\boldsymbol{\mathbf{p}}}^{*}_{j}}{2})\}$ (14)
A property of voronoi diagrams is that, for each object $SQ_{o}$, the intersection of all the spaces created by the bisector hyperplanes that contain the object forms the polyhedral cell corresponding to that object. The collection of all such polyhedral cells constitutes the complete Voronoi diagram. Initially, We define a bounding box (world space $W$) to appropriately represent the workspace. Consequently, the region $V_{i}$ corresponding to each cluster is represented as the intersection of all the hyperplanes that define its polyhedral cell within this bounding box.
$V_{i}=\bigcup_{j=0}^{m}W\cap HP_{i,j}.$ (15)
The result of Eq. 15 is shown as Fig.2(d). We observe most obstacles belong to one region, while there are two overlap obstacles belongs to same region. The steps to generate a Voronoi diagram with cluster is shown in Alg. 2.
Algorithm 2 Generate Voronoi diagram
1: Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$.
2: Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. procedure Create Voronoi(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$) Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. procedure Create Voronoi(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$)
3:      Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. procedure Create Voronoi(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$) Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. procedure Create Voronoi(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$) Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. procedure Create Voronoi(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$) for $i=1,\dots,K$ do Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. procedure Create Voronoi(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$) Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. procedure Create Voronoi(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$) for $i=1,\dots,K$ do
4:          Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. procedure Create Voronoi(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$) Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. Cluster $C={[C_{1},\dots,C_{K}]}$, the world space $W$. procedure Create Voronoi(${\boldsymbol{\mathbf{q}}}_{\text{near}},\delta{\boldsymbol{\mathbf{u}}},\epsilon$) for $i=1,\dots,K$ do $C_{i}=C(i)$ $C_{i}=C(i)$
5:          $C_{i}=C(i)$ $C_{i}=C(i)$ $C_{i}=C(i)$ for $j=i,\dots,K$ do $C_{i}=C(i)$ $C_{i}=C(i)$ for $j=i,\dots,K$ do
6:               $C_{i}=C(i)$ $C_{i}=C(i)$ for $j=i,\dots,K$ do $C_{j}=C(j)$ $C_{j}=C(j)$
7:               $C_{j}=C(j)$ $C_{j}=C(j)$ $C_{j}=C(j)$ if $i\neq j$ then $C_{j}=C(j)$ $C_{j}=C(j)$ if $i\neq j$ then
8:                   $C_{j}=C(j)$ $C_{j}=C(j)$ if $i\neq j$ then $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{{\boldsymbol{\mathbf{p}}}_{i},{\boldsymbol{\mathbf{p}}}_{j}}{\arg\min}\;d({\boldsymbol{\mathbf{p}}}_{i},{\boldsymbol{\mathbf{p}}}_{j})$, where ${\boldsymbol{\mathbf{p}}}_{i}\in SQ^{M}_{o,i}\in C_{i},\;{\boldsymbol{\mathbf{p}}}_{j}\in SQ^{M}_{o,j}\in C_{j}$ $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{{\boldsymbol{\mathbf{p}}}_{i},{\boldsymbol{\mathbf{p}}}_{j}}{\arg\min}\;d({\boldsymbol{\mathbf{p}}}_{i},{\boldsymbol{\mathbf{p}}}_{j})$, where ${\boldsymbol{\mathbf{p}}}_{i}\in SQ^{M}_{o,i}\in C_{i},\;{\boldsymbol{\mathbf{p}}}_{j}\in SQ^{M}_{o,j}\in C_{j}$ $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{{\boldsymbol{\mathbf{p}}}_{i},{\boldsymbol{\mathbf{p}}}_{j}}{\arg\min}\;d({\boldsymbol{\mathbf{p}}}_{i},{\boldsymbol{\mathbf{p}}}_{j})$, where ${\boldsymbol{\mathbf{p}}}_{i}\in SQ^{M}_{o,i}\in C_{i},\;{\boldsymbol{\mathbf{p}}}_{j}\in SQ^{M}_{o,j}\in C_{j}$
9:               $(p_{i}^{*},p_{j}^{*})\leftarrow\underset{{\boldsymbol{\mathbf{p}}}_{i},{\boldsymbol{\mathbf{p}}}_{j}}{\arg\min}\;d({\boldsymbol{\mathbf{p}}}_{i},{\boldsymbol{\mathbf{p}}}_{j})$, where ${\boldsymbol{\mathbf{p}}}_{i}\in SQ^{M}_{o,i}\in C_{i},\;{\boldsymbol{\mathbf{p}}}_{j}\in SQ^{M}_{o,j}\in C_{j}$ $HP_{i,j}\leftarrow$ Eq. 14. $HP_{i,j}\leftarrow$ Eq. 14.
10:               $HP_{i,j}\leftarrow$ Eq. 14. $V_{i}=\bigcup_{j=0}^{m}W\cap HP_{i,j}$ Eq. 15. $V_{i}=\bigcup_{j=0}^{m}W\cap HP_{i,j}$ Eq. 15. $V_{i}=\bigcup_{j=0}^{m}W\cap HP_{i,j}$ Eq. 15.
11:      $V_{i}=\bigcup_{j=0}^{m}W\cap HP_{i,j}$ Eq. 15. return $V$ return $V$ return $V$
III-C Generate graph via Voronoi diagram
We create a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where each $V_{i}$ corresponds to a polygon with multiple vertices and edges. Each vertex of the polygon is represented as a node $v_{i}$ in a graph $\mathcal{V}$, where nodes are connected by edge on polygon $\mathcal{E}$. The weight assigned to each edge corresponds to the Euclidean distance between the connected nodes.
We notice that there are certain regions (holes) in the diagram that do not belong to any cell $V_{i}$. This occurs due to the linear approximation used in constructing the diagram, but it has few impact since our primary objective is to identify safe paths along the roadmap, which we obtain using this straightforward approach. The same methodology can be extended to 3D, where we obtain polyhedra instead of polygons.
Meanwhile, to address these holes, we connect nodes where the distance between them is smaller than a predefined threshold $h$, which is determined based on the overall size of the workspace, ensuring that the connections adequately represent feasible paths through the environment. In other words, we adjust $h$ relative to the scale of the world to maintain a practical and navigable graph structure.
$\mathcal{E}$ =$=\mathcal{E}\cup\{v_{i},v_{j}\},\;\forall v_{i},v_{j}\in V,\norm{v_{i}-v_{j}}\leq h$ (16)
Given any initial pose and target pose, we first project these poses onto their respective closest edges $\mathcal{E}$. Once projected, we determine the shortest path through the graph by classical graph planning algorithm (e.g., Dijkstra) to find solution path via nodes $\mathcal{V}_{sol}=[v_{1},v_{2},\dots,v_{i}]^{T}$.
III-D Planning with orientation
In subsection A, we determined the longest axis of the robot $SQ_{r}$. To control the robot’s orientation while navigating through narrow passages, the robot should move along the direction of its longest axis. This direction is computed based on the nodes in the path solution, $\mathcal{V}_{sol}$, which provides the optimal orientation for each step of the path.
III-D1 2D Case
In the 2D case, we plan the robot’s motion using ${\boldsymbol{\mathbf{T}}}({\boldsymbol{\mathbf{p}}},{\boldsymbol{\mathbf{\theta}}})\in SE(2)$, where ${\boldsymbol{\mathbf{\theta}}}\in\mathfrak{so}(2)$. For each edge in the path, the direction is determined by the orientation of the robot’s longest axis relative to the edge,
$\Delta v$ $:=[\Delta v_{x},\Delta v_{y}]^{T}=v_{i+1}-v_{i}\in\mathbb{R}^{2}$ (17)
${\boldsymbol{\mathbf{\theta}}}$ $=\arctan(\Delta v_{y},\Delta v_{x})$ (18)
III-D2 3D
In 3D case, we are planning for ${\boldsymbol{\mathbf{T}}}({\boldsymbol{\mathbf{p}}},{\boldsymbol{\mathbf{\theta}}})\in SE(3)$, where ${\boldsymbol{\mathbf{\theta}}}\in\mathfrak{so}(3)$. Since we are navigating through nodes located on the surfaces of polyhedra, these surfaces provide the maximum clearance orientation for the robot. To ensure efficient navigation, the robot’s ($SQ_{r}$) longest axis is aligned forward along the direction of travel, while the shortest axis is oriented perpendicular (normal) to the halfplane $HP_{i,j}$. This configuration maximizes the robot’s ability to fit through narrow passages and maintain optimal clearance. The rotation vector required to adjust the robot’s orientation is computed based on the vertices of the surfaces, ensuring that the robot aligns itself as it moves from node to node on the surface.
${\boldsymbol{\mathbf{r}}}_{1}$ $=\hat{n}(v_{i+1}-v_{i})$
${\boldsymbol{\mathbf{r}}}_{3}$ $={\boldsymbol{\mathbf{n}}}_{i,j},\;v_{i+1},v_{i}\in HP_{i,j}$
${\boldsymbol{\mathbf{r}}}_{2}$ $={\boldsymbol{\mathbf{r}}}_{3}\times{\boldsymbol{\mathbf{r}}}_{1}$
${\boldsymbol{\mathbf{R}}}$ $=[{\boldsymbol{\mathbf{r}}}_{1},{\boldsymbol{\mathbf{r}}}_{2},{\boldsymbol{\mathbf{r}}}_{3}]$ (19)
After determining the rotation matrix at each edge, we compute the corresponding Lie algebra ${\boldsymbol{\mathbf{\theta}}}\in\mathfrak{so}(3)$. This is done by applying the matrix logarithm $\log$ to the rotation matrix, which gives a skew-symmetric matrix, then $\vee$ represents extract skew-symmetric matrix to vector.
${\boldsymbol{\mathbf{\theta}}}$ $=\log_{\vee}{\boldsymbol{\mathbf{R}}}$ (20)
III-E Smooth Interpolation Using DMP
Now that we have the orientation at each edge and the nodes as via points, we need a smooth trajectory for the robot to follow. First, we interpolate all the nodes using polynomials to obtain the trajectory ${\boldsymbol{\mathbf{p}}}(t)$. However, this trajectory may still appear zigzagged, and it is not necessary to pass through every node, especially in areas with many redundant nodes (e.g., holes in the diagram). To address this, we employ Dynamical Movement Primitives (DMP) to learn a smooth version of this trajectory. DMP is a technique that parameterizes and smooths trajectories [21]. Classical DMP can be used to map a finite dimensional set of parameters ${\boldsymbol{\mathbf{W}}}\in\mathbb{R}^{K\times P}$ where $P$ is the number of Radial Basis Functions (RBFs) per degree-of-freedom (DOF) into smooth and differentiable functions.
DMP $:({\boldsymbol{\mathbf{W}}},{\boldsymbol{\mathbf{u}}}_{start},{\boldsymbol{\mathbf{u}}}_{goal},T)\mapsto{\boldsymbol{\mathbf{u}}}_{{\boldsymbol{\mathbf{W}}}}(t),\quad t\in[0,T].$ (21)
where ${\boldsymbol{\mathbf{u}}}_{{\boldsymbol{\mathbf{W}}}}(t)$ represents the desired pose. We first augment position and lie algebra, so we have dataset ${\boldsymbol{\mathbf{D}}}=[{\boldsymbol{\mathbf{p}}}(t),{\boldsymbol{\mathbf{\theta}}}(t)],t\in[0,T]$, then we apply locally weighted regression (LWR) to learn parameter ${\boldsymbol{\mathbf{W}}}$ from dataset, which is used to generate the new trajectory.
${\boldsymbol{\mathbf{W}}}$ $=\textsc{LWR}({\boldsymbol{\mathbf{D}}})$ (22)
The resulting trajectory is shown in Fig. 2(e), where the blue line represents the original edges of the solution path, and the pink trajectory generated by DMP shows the smoothed path.
IV Experiment: Planar retrieval in clutter
Retrieving a target object from clutter is a common task in daily life (e.g., from shelves). However, due to height restrictions, retrieving objects from the top is not feasible, making it challenging to find a collision-free path for retrieval [22]. We designed a toy experiment to simulate a similar scenario and test the performance of our algorithm in a 2D environment.
IV-A Experiment design and results
Figure 3: Process for executing the experiment. Step 1: The Kinova robot scans a table crowded with objects. Step 2: Objects are segmented, and the target object is selected as $SQ_{r}$. Step 3: The remaining obstacles $SQ_{o}$ are expanded, where the impossible passage are eliminated. Step 4: Trajectory planning is performed with the desired orientation using DMP.
The process of this experiment is illustrated in Fig. 3. We set up a table with several objects, some of which are placed to create narrow passages that too difficult to pass through. A Kinova robot equipped with a depth camera is positioned on the table, and the camera is calibrated. The robot raises to its highest position while keeping its gaze downwards, and images are captured.
We segment the objects using the SAM toolbox [23], from which we obtain point clouds representing the boundaries of all obstacles. Superellipse fitting is then applied to these point clouds to obtain the obstacle representation [17]. Next, we applied our framework with the help of bensolve toolbox [24], and the results are shown at Fig. 3. We manually select an object $SQ_{r}$ for retrieval. The robot moves to ${\boldsymbol{\mathbf{p}}}_{r},{\boldsymbol{\mathbf{\theta}}}_{r}$ to grasp the object. Subsequently in the right bottom figure, the size of other obstacles are expanded by $a_{r,1}$. After expansion, objects close to each other overlap, indicating that they belong to the same region $V_{i}$. Planning with orientation ensures that the long axis of the object aligns with the edge in Voronoi diagram. The final trajectory is sent to the robot, which executes the trajectory, allowing the object to safely pass through the cluttered environment.
We set up two types of tables for the experiment. In one scenario, the density of obstacles on the table is relatively moderate. In the other, the arrangement is denser, with passages between obstacles that may be narrower than the shortest length of the object (as shown in Fig.3). Each scenario was tested five times with different starting points and target positions, and all trials were successful.
TABLE I: Benchmark for 2D scenario.
Metric A* RRT* Genetic Algorithm Voronoi - planning
computing time (s) 0.42 3.16 5.78 0.77
arc length (m) 0.41 0.76 0.45 0.61
min-distance (mm) 0.7 7.1 1.1 14.2
solution [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
IV-B Benchmark for 2D scenario
All the methods compared here are implemented in MATLAB, for better comparison. We have chosen the same start and end points in all the cases, to visually see the differences between the optimal paths generated by each algorithm. A summary of the comparison is shown in the table I.
In this comparison, we evaluate three key metrics: computation time, arc length, and minimum distance. All final trajectories are plotted after each algorithm’s execution. The computation time includes all phases of the algorithm, the arc length represents the total length of the planned trajectory, and the minimum distance refers to the closest distance between the robot $SQ_{r}$ to any obstacle $SQ_{o}$.
We observe that our algorithm completes within 1 second in this scenario (though the computation time depends on the complexity of the environment and the programming language used). The key strength of our approach is safety, as the robot always passes through the bisector between obstacles, maximizing clearance. This minimum distance is particularly important in cases involving small objects, as they are harder to control accurately, increasing the likelihood of collisions. This trend is evident from the figures in Table I. Algorithms like A*, RRT* and Genetic Algorithm often generate trajectories that come very close to obstacles, increasing the risk of collisions. On the other hand, while our algorithm consistently maintains a safer distance, it tends to produce longer trajectories, which is one disadvantage.
V Simulation: drone navigation through narrow passage
In this section, we apply our algorithm into 3D world. A suitable example is drone flying in a crowded world. Assume all the obstacles can be approximate by SQ, including the drone, which looks like a ellipsoid.
V-A Experiment setup
In this experiment, we designed three different maps: moderate, dense, and narrow. Moderate indicates a medium density of obstacles in the 3D environment, with dense and narrow reflecting higher obstacle density. We first generate the map and generate a point cloud as input for GCOPTER, where we test their algorithm by their open source code [20]. We randomly selected start and target points five times in each map to make comparisons.
(a) Our method v.s. GCOPTER (open-source) generate similar trajectory in a moderate crowd space.
(b) Our method v.s. GCOPTER (open-source) in a comparable crowd space, our method prefer safer path.
Figure 4: Our method v.s. GCOPTER (open-source) in a moderate and dense environment.
Fig. 4 illustrates the moderate and dense maps. In these figures, GCOPTER’s trajectory is shown in blue, while our method’s trajectory is depicted in green. In Fig.4(a), where the space is not highly crowded, the trajectories generated by both approaches are similar, though our method shows slight zigzagging compared to GCOPTER. However, as the environment becomes more crowded, our method tends to prioritize safety by maintaining a greater distance from obstacles, as seen in Fig.4(b).
In the third scenario, we created a crowded passage environment, a setup commonly used in real-world applications [4]. In this case, we manually positioned four SQs as pillars, leaving only one passage wide enough for the drone to pass through. Specifically, the gap between the obstacles is smaller than the drone’s long axis but larger than its short axis, allowing it to pass through the narrow passage. The other two passages are too narrow for the drone to pass.
The results are shown in Fig. 5, our method successfully identifies and eliminates the gaps corresponding to the two impassable passages and finds the viable path through the narrow passage, as demonstrated in Fig. 5(a). Additionally, the drone tilts at an angle to navigate through the passage, as shown in the zoomed-in view in Fig. 5(b). In this figure, the blue ellipsoid represents the drone, while the transparent shapes depict the surrounding obstacles. On the other hand, we tested the open-source GCOPTER, which did not achieve similar results in the same environment. While the GCOPTER paper references capabilities for navigating narrow spaces, the open-source implementation lacks this specific functionality, which may contribute to the observed performance difference.
(a) A narrow passage in a crowd world making hard for planning.
(b) the drone performs a tilt angle to pass through the passage.
Figure 5: Our method v.s. GCOPTER in a passage environment.
In Table II, we record the average outcomes across various tests and scenarios. We list planning time [^1] Planning time is measured from when the start and target points are provided until the final trajectory is generated. The time spent calculating the closest points between each SQ and generating the corresponding polygons are not included in this measurement., min-distance between drone to the closest obstacle, and arc length of the final trajectory. The results in the table reflect the same conclusions as the figures and analysis discussed earlier. Both methods show similar planning times. However, the average arc length is longer in our method, as it prioritizes finding a safer path, which is evident from the greater minimum distance from obstacles.
TABLE II: Our method v.s. GCOPTER.
method moderate dense narrow
planning time(ms) GCOPTER (open-source) 31.56 33.86
proposed 35.76 53.90 64.60
min-distance(m) GCOPTER (open-source) 1.28 0.87
proposed 1.71 1.46 0.28
arc length(m) GCOPTER (open-source) 27.06 24.03
proposed 40.01 32.02 33.47
VI Conclusion
This study presents an effective approach that integrates SQ with Voronoi diagrams for maximum clearance path planning. By representation of SQs provides versatile parametrization of complex shapes, we constructed a Voronoi diagram roadmap to generate a sparse graph of nodes and edges, significantly reducing computational complexity while ensuring safe and efficient paths. Additionally, the bisectors naturally support the robot’s orientation during navigation.
Our method is applicable to both 2D and 3D cases, and is capable of distinguishing between narrow passages and impassable passages. We demonstrated the effectiveness of our algorithm in 2D with a retrieval example. Furthermore, we compared our approach with a cutting edge drone planner, showing consistent results with those from the 2D example.
However, our method has limitations when dealing with non-convex obstacles, as it assumes all obstacles can be approximated by convex SQs. Additionally, the toolbox we used, Bensolve, is limited to convex polygons, meaning the polygonal representation of each cluster is also convex.
In conclusion, the integration of Voronoi diagrams with superquadric representations offers a promising approach for enhancing path planning capabilities in robotics and autonomous systems, particularly in environments that require precise and safe navigation. In the future, we aim to extend the planner to include robotic manipulators, allowing for both end-effector and manipulator path planning to achieve self-collision avoidance.